# A Frame-based Attention Interpretation Method for Relevant Acoustic Feature Extraction in Long Speech Depression Detection

### v1

@@ Abstract
Speech-based depression detection tools could help early screening of depression. Here, we address two issues that may hinder the clinical practicality of such tools: segment-level labelling noise and a lack of model interpretability. We propose a speech-level Audio Spectrogram Transformer to avoid segment-level labelling. We observe that the proposed model significantly outperforms a segment-level model, providing evidence for the presence of segment-level labelling noise in audio modality and the advantage of longer-duration speech analysis for depression detection. We introduce a frame-based attention interpretation method to extract acoustic features from prediction-relevant waveform signals for interpretation by clinicians. Through interpretation, we observe that the proposed model identifies reduced loudness and F0 as relevant signals of depression, which aligns with the speech characteristics of depressed patients documented in clinical studies.
@@
frame-based attention interpretation, speech analysis, depression detection
@@ 1 Introduction
Depression is a common mental disorder, characterised by prolonged low mood, and loss of interest in activities, with an estimated 5% of adults suffering from it globally [1]. Recently, more research attention has been placed on developing automatic depression screening tools, using Deep Neural Networks (DNNs) to analyse patients’ speech. These tools have the potential to monitor individuals’ risk of depression at early stages and assist clinicians in providing rapid interventions.
This paper addresses two issues that may hinder the clinical implementation of speech-based screening tools for depression: noise that could result from segment-level labelling and a lack of model interpretability.
@@
First, segmentation of data sequences is a common approach to avoid processing long sequences for training DNNs. For instance, [2] and [3] segmented audio data sequences along the temporal dimension, each labelled according to the participant’s overall label. However, assuming that there are speech segments from depressed patients that contain no depression-relevant information (i.e. if depressive markers are not evenly spread across the whole speech), and are nonetheless labelled “depressed”, this segment-level labelling approach can add noise to the model training and consequently reduce prediction accuracy in clinical practice.
@@
To avoid labelling noise in social media analysis, [4] applied a post-level encoder to first encode social media posts from the same user into fixed-size embeddings. The sequence of embeddings from each user is then fed to a user-level encoder for a user-level depression classification given a single label. While segment-level labelling noise may be obvious in social media analysis (i.e., text modality), where some posts do not explicitly contain depression-related utterances, it is not obvious for speech analysis (i.e., audio modality).
Inspired by [4], our model performs sentence-level segmentation for long speech analysis, whereby speech is first segmented into audio segments consisting of natural sentences. We then apply a sentence-level encoder to encode the segments into a sequence of embeddings, which can then be processed by a speech-level encoder to make a single prediction for each participant, dispensing with the need for segment-level labelling. We then compare the model’s performance to a model that learns at a segment level (segment-level model) to investigate segment-level labelling noise in speech-based depression detection.
@@
Second, predictions made by DNNs are hardly interpretable. The lack of interpretability can hinder the application of speech-based depression screening tools in clinical practice [5, 6]. Attention scores have been suggested as a means of model interpretation because they provide understandable weight distributions over input features [7]. In their design of a Hierarchical Attention Network (HAN), for instance, [8] acquired attention scores from attention layers to infer the importance of tokens in social media tweets for depression detection. Intuitively, input tokens at the sequence positions with high attention scores contribute more to depression detection.
@@
However, this attention-based interpretation approach is insufficient for two reasons: firstly, it ignores the computations that happened before the attention layers [9]. Specifically, in [8], before attention operations, each token has already been contextualised by every other token by a bidirectional Gated Recurrent Unit (biGRU). Therefore, representations produced from the biGRU at each input position capture not only the information of the token itself but also the context of the whole tweet. Consequently, the attention scores at later layers do not map directly onto the input tokens, but rather onto their context-enriched representations, making it insufficient to interpret the attention scores as representing the importance of individual tokens.
Secondly, it has been demonstrated that raw attention weights often do not correlate with gradient-based measures of feature significance [10], which questions the validity of using raw attention weights for model interpretation. To address these insufficiencies, we design the proposed model to be based entirely on the attention mechanism so that we can track interactions between input features at every attention layer using gradient-weighted attention maps [9]. By tracking interactions, we introduce a novel frame-based attention interpretation method to highlight Mel-spectrogram frames relevant to depression detection, with which specific acoustic features can be extracted for interpretation by human experts.
@@ 2.1 Data pre-processing
2.2 Speech-level model architecture
The proposed speech-level model is based entirely on the attention mechanism [13]. First, the model consists of a sentence-level processing block, which encodes sentence-level spectrograms using a pre-trained frame-based Audio Spectrogram Transformer (AST) [14]. Instead of partitioning the spectrogram into patches of size 16 × 16, the frame-based AST splits the spectrogram into frames of size 128 × 2 along the temporal dimension. We add attention masks to ignore padding. After the processing, the [cls] token from the audio data then represents the sentence-level audio data.
@@ 2.3 Frame-based attention interpretation
Figure 2: Violin plot of relevant acoustic feature distributions between true positives (n = 60) and true negatives (n = 41). Relevant acoustic features were extracted from the waveform signals temporally corresponding to the spectrogram frames with relevancy scores higher than 0.3 from the five most relevant sentences of each sample. For true negatives, gradients of the output for the “normal” class with respect to the attention scores were used to weigh the attention maps. Statistical significance was assessed using the Mann-Whitney U test, with Bonferroni correction applied for multiple comparisons. Significance levels were denoted as: $ns$ for not significant; * for $p\leq 0.002$, indicating significant differences; ** for $p\leq 0.0004$; and *** for $p\leq 4\times 10^{-5}$.
@@ 3.3 Implementation details
For both the speech-level and segment-level models, we froze the first nine layers and randomly reinitialised the last three layers of the pre-trained AST. Both models were trained using an Adam optimizer [16] with a learning rate of $3\times 10^{-5}$ and a weight decay of 0.01. For a fair comparison, training was stopped at the onset of overfitting (epoch 9 for the speech-level model and epoch 8 for the segment-level model) as indicated by the cessation of further decrease in development loss. While we acknowledge that this approach will lead to overly optimistic estimates of performance for both models, we note that this work does not aim to benchmark performance on the D-Vlog dataset.
@@ 3.4 Model performance
Figure 3: Perturbation test results for the speech-level model. Accuracy was computed using a decision threshold of 0.527, reflecting the prevalence of depression in the data. In the first test, sentence-level spectrograms were incrementally excluded in descending order of relevance, 10 at a time. In the second test, only frames with a relevance score above 0.3 within these sentence spectrograms were removed. The results of these two tests were benchmarked against random exclusions: one where sentence spectrograms were randomly excluded, and another where random frames—amounting to 30% within any given relevant sentence spectrogram—were removed. Note that the accuracy never drops below 47.18% for all tests, likely because the model defaults to positive predictions in the absence of input signals due to the slight class imbalance in the training set.
@@ 3.5 Perturbation study
We conducted perturbation tests to explore the impact of omitting relevant sentence-level spectrograms or frames on our speech-level model’s performance. As shown in Figure 3, with detailed test descriptions provided in the figure caption, the selective removal of relevant frames led to a marked decline in model accuracy, substantially outpacing the accuracy degradation observed with random frame removal. The selective exclusion of relevant sentence-level spectrograms resulted in a further reduction in accuracy, which is marginally more pronounced, especially for the first 10 sentences, than that observed with random removal of sentences. This observation implies differences between sentence-level spectrograms, or speech segments, in terms of signal importance for model predictions.
@@ 3.6 Relevant acoustic feature extraction
For the final step of the interpretation method, we locate the waveform signals temporally corresponding to the relevant spectrogram frames and extract a set of acoustics features from these signals using OpenSMILE (v2.5.0). Figure 2 demonstrates the results of interpreting the proposed model by comparing the distribution differences in terms of the eGeMAPS (v02) [19] feature set between true positive/negative samples (see a detailed procedure description in the figure caption). We observe that the proposed model, for instance, identifies reduced loudness and fundamental frequency (F0) as relevant signals of depression, which aligns with the speech characteristics of depressed patients documented in clinical studies [20].
@@ 4 Conclusions
This work provides evidence for the presence of segment-level labelling noise in the audio modality for depression detection, showing the advantage of longer-duration speech analysis for clinical predictions over those based on short speech segments. Crucially, this work demonstrates the feasibility of audio interpretation for DNNs-based disease detection, where relevant acoustic features can be extracted for human interpretation. For future work, we propose that the interpretation method can be applied to systematically study the behaviour of speech-based disease detection systems. For instance, a system’s characteristics (e.g., reliance on certain acoustic features) in making false positives/negatives can be documented prior to their clinical implementations, which allows clinicians to critically assess the reliability of certain model predictions.

### v3

@@ Abstract
Speech-based depression detection tools could aid early screening. Here, we propose an interpretable speech foundation model approach to enhance the clinical applicability of such tools. We introduce a speech-level Audio Spectrogram Transformer (AST) to detect depression using long-duration speech instead of short segments, along with a novel interpretation method that reveals prediction-relevant acoustic features for clinician interpretation. Our experiments show the proposed model outperforms a segment-level AST, highlighting the impact of segment-level labelling noise and the advantage of leveraging longer speech duration for more reliable depression detection. Through interpretation, we observe our model identifies reduced loudness and F0 as relevant depression signals, aligning with documented clinical findings. This interpretability supports a responsible AI approach for speech-based depression detection, rendering such tools more clinically applicable.
@@
depression detection, speech foundation model, responsible AI, audio interpretation
@@ 1 Introduction
Depression is a common mental disorder, characterised by prolonged low mood, and loss of interest in activities, with an estimated 5% of adults suffering from it globally [1]. Recently, more research attention has been placed on developing automatic depression screening tools, using Deep Neural Networks (DNNs) to analyse patients’ speech. These tools have the potential to monitor individuals’ risk of depression at early stages and assist clinicians in providing rapid interventions. This paper addresses two issues that may hinder the clinical applicability of such clinical tools: predictions relying on short speech segments and a lack of model interpretability.
@@
First, segmentation of data sequences is a common approach to avoid processing long sequences for DNN-based depression detection. For instance, [2] and [3] segmented audio data sequences along the temporal dimension, each labelled according to the subject’s overall diagnosis. However, assuming that there are speech segments from depressed patients that contain no depression-relevant information (i.e. if depressive markers are not evenly spread across the whole speech), and are nonetheless labelled “depressed”, this segment-level labelling approach can add noise to the model training and consequently reduce prediction accuracy in clinical practice.
@@
To avoid labelling noise in social media analysis, [4] applied a post-level encoder to first encode social media posts from the same user into fixed-size embeddings. The sequence of embeddings from each user is then fed to a user-level encoder for a user-level depression classification given a single label. While segment-level labelling noise may be obvious in social media analysis (i.e., text modality), where some posts do not explicitly contain depression-related utterances, its impact on speech-based depression detection (i.e., audio modality) demands further investigation.
@@
Second, speech-based depression detection tools are hardly interpretable, hindering their clinical implementations in practice [5, 6]. In text-based depression detection (e.g., social media analysis), attention scores from transformer models have been used for model interpretation because they provide understandable weight distributions over input features [7]. For instance, [8] acquired attention scores from their Hierarchical Attention Network (HAN) to infer the importance of tokens in social media tweets for depression detection. Intuitively, input tokens at the sequence positions with high attention scores contribute more to depression detection.
@@
However, this interpretation approach is insufficient because it ignores the computations that happened before the attention layers [9]. Specifically, in [8], before attention operations, each token has already been contextualised by every other token by a bidirectional Gated Recurrent Unit (biGRU). Therefore, attention scores derived from later layers do not map directly onto the input tokens, but rather onto their context-enriched representations, rendering the interpretation less precise. Additionally, raw attention scores often poorly correlate with gradient-based feature significance [10], questioning their validity for model interpretation. More importantly, this attention-based interpretation approach is not readily applicable to speech-based depression detection because unlike text, which is directly understandable to humans, raw audio signals require further processing to be human-interpretable.
To address these insufficiencies, we design an interpretable speech foundation model approach, leveraging pre-trained Audio Spectrogram Transformer (AST) [11], to make end-to-end depression detection using one’s long speech instead of short segments. We compare the model to a segment-level AST to highlight the impact of segment-level labelling noise and the advantage of using longer speech duration for more reliable depression detection. Importantly, based on gradient-weighted attention maps [9], we introduce a novel frame-based interpretation method to extract human-interpretable acoustic features relevant to depression detection, making such speech foundation models interpretable for clinicians, and therefore more clinically applicable, for the first time.
@@ 2.1 Data pre-processing
2.2 Speech-level Audio Spectrogram Transformer
The proposed speech foundation model approach is based entirely on the attention mechanism [14]. First, the model consists of a sentence-level processing block, which encodes sentence-level spectrograms using a pre-trained frame-based AST. Instead of partitioning the spectrogram into patches of size 16 × 16, the frame-based AST splits the spectrogram into frames of size 128 × 2 along the temporal dimension. We add attention masks to ignore padding. After the processing, the [cls] token from the audio data then represents the sentence-level audio data.
@@ 2.3 Frame-based attention interpretation
Figure 2: Violin plot of relevant acoustic feature value distributions (residualized for sex and standardized) between true positives (n = 60) and true negatives (n = 41). Relevant acoustic features were extracted from the waveform signals temporally corresponding to the spectrogram frames with relevancy scores higher than 0.3 from each sample’s five most relevant sentences. For true negatives, gradients of the output for the “normal” class regarding the attention scores were used to weigh the attention maps. Statistical significance was assessed using the Mann-Whitney U test, with Bonferroni correction applied for multiple comparisons. Significance levels were denoted as: $ns$ for not significant; * for $p\leq 0.002$, indicating significant differences; ** for $p\leq 0.0004$; and *** for $p\leq 4\times 10^{-5}$.
@@ 3.3 Implementation details
For both the speech-level and segment-level models, we froze the first nine layers and randomly reinitialised the last three layers of the pre-trained AST. Both models were trained using an Adam optimizer [16] with a learning rate of $3\times 10^{-5}$ and a weight decay of 0.01. For a fair comparison, training was stopped at the onset of overfitting (epoch 9 for the speech-level model and epoch 8 for the segment-level model) as indicated by the cessation of further decrease in development loss. While we acknowledge that this approach may lead to overly optimistic estimates of performance for both models (lacking an independent test set due to limited sample size), we note that this work does not aim to benchmark performance on the D-Vlog dataset. Instead, we aim to provide robust evidence for the advantage of using a patient’s long speech over short segments for more reliable depression detection in clinical practice.
@@ 3.4 Model performance
Figure 3: Perturbation test results for the proposed model. Accuracy was computed using a decision threshold of 0.527, reflecting the depression prevalence in the data. In the first test, sentence-level spectrograms were incrementally excluded in descending order of relevance, 10 at a time. In the second, only frames with a relevance score above 0.3 within these sentences were removed. The results of these two tests were benchmarked against random exclusions: one with random sentence spectrogram exclusions, and another with random frames exclusions (30% within relevant sentence spectrogrames). Accuracy never drops below 47.18%, likely due to the model’s bias towards positive predictions in the presence of a slight class imbalance.
@@ 3.5 Perturbation study
We conducted perturbation tests to explore the impact of omitting relevant sentence-level spectrograms or frames on our speech-level model’s performance. As shown in Figure 3, with detailed test descriptions provided in the figure caption, the selective removal of relevant frames led to a marked decline in model accuracy, substantially outpacing the accuracy degradation observed with random frame removal. The selective exclusion of relevant sentence-level spectrograms resulted in a further reduction in accuracy, which is marginally more pronounced, especially for the first 10 sentences, than that observed with random removal of sentences. This observation implies differences between sentence-level spectrograms, or speech segments, in terms of signal importance for depression detection.
@@ 3.6 Relevant acoustic feature extraction
For the final step of the interpretation method, we locate the waveform signals temporally corresponding to the relevant spectrogram frames and extract a set of acoustics features from these signals using OpenSMILE (v2.5.0). We used the eGeMAPS (v02) feature set, which was designed to index affective changes in voice [19]. Figure 2 demonstrates the distribution differences in relevant acoustic feature values between true positive/negative samples. For a preliminary interpretation, we note that the model identifies reductions in loudness and fundamental frequency (F0) as relevant signals of depression, which aligns with the speech characteristics of depressed patients documented in clinical studies [20]. Future work is planned to validate the clinical relevance of our method using clinical speech datasets of depression (e.g., DAIC-WOZ [21]).
@@ 4 Conclusions
This work presents an interpretable speech foundation model approach that leverages pre-trained parameters, enabling scalability with more data, and uses patients’ long speech for more reliable depression detection in clinical practice. Crucially, it advances responsible AI by introducing a feasible audio interpretation method that reveals prediction-relevant acoustic features for clinician interpretation, which enhances early screening tools’ clinical applicability. With such interpretability, future works may systematically study the behaviour of speech-based depression detection systems. For instance, a system’s characteristics (e.g., reliance on certain acoustic features) in making false positives/negatives can be documented prior to their clinical implementations, which allows clinicians to critically assess the reliability of certain model predictions.